OpenStudy (anonymous):

volume of solid bounded by cone of x^2+y^2=z^2 and the paraboloid of x^2+y^2=z ? and someone draw the graph and explain thanks....use triple integral

OpenStudy (anonymous):

wow i can now visualize it clearly @cinar but now for the triple integral formula i dun know the region of integration lies could u explain to me and is it cone minus paraboloid or vise versa? ty....

OpenStudy (anonymous):

examine example 5 http://www.math24.net/calculation-of-volumes-using-triple-integrals.html

OpenStudy (anonymous):

wow kk steps are explained clearly but do u know which minus which?? cone minus paraboloid or paraboloid minus cone that confusing me only .....

OpenStudy (anonymous):

choose one of them we know that volume can not be negative, if it is negative then it is versa.. I guess second one has larger volume, you can see it at graph

OpenStudy (anonymous):

I forgot these stuff sorry..

OpenStudy (turingtest):

x^2+y^2=z^2 means that z=sqrt(x^2+y^2) x^2+y^2=z means that z=x^2+y^2 so z will have a greater value in the cone than the parabaloid, hence it is on top

OpenStudy (anonymous):

you are saying opposite of what you thinking

OpenStudy (turingtest):

you are right, I meant it has a greater value in the parabaloid, sorry

OpenStudy (anonymous):

hmm means its paraboloid minus cone? @TuringTest

OpenStudy (turingtest):

yes, because\[\sqrt{x^2+y^2}\le x^2+y^2\]anyway, this should be easier in cylindrical coordinates \[x^2+y^2=r^2\] I have to go take a Spanish test, good luck, I'll check on this when I return :D

OpenStudy (anonymous):

if i do both paraboloid minus cone and vise versa will i get exact value but diff sign one positive one negative? @TuringTest

OpenStudy (anonymous):

http://www.physicsforums.com/showthread.php?t=15266

OpenStudy (turingtest):

yes I think you could change the sign at the end we need the intersection of these two surfaces converting this to cylindrical coordinates\[x^2+y^2=r^2=z^2\to z=\pm r\]\[x^2+y^2=r^2=z\]setting the z of both equations equal we get\[r=r^2\to r=0,1\]hence our bounds are\[r\le z\le r^2\]\[0\le r\le1\]\[0\le\theta\le2\pi\]and you can now integrate